hello! my name is daniela and i dont understand how to figure out
this question can u please help me? like i dont understand what steps to do (if that helps)
what is the range of the fuction f(x)=4x-3 when the domain is -1,0,3?
thank you very much :)

Answer by stanbon(48535)

what is the range of the fuction f(x)=4x-3 when the domain is -1,0,3?
f(-1)=4(-1)-3=-7
f(0)=4(0)-3=-3
f(3)=4(3)-3=9
The range is {-7,-3, 0}
Cheers,
Stan H.